OpenStudy (anonymous):

there are 785 students in the senior class. if there are 77 more females in the class than males, how many male and female seniors are there in the class?

OpenStudy (therealmeeeee):

let me see

OpenStudy (johnweldon1993):

This is a quick substitution problem... 785 students in all...male and female....so m + f = 785 There are 77 more females...so we have f = m + 77 Now that we know what 'f' equals...we can substitute this into the original problem... m + f = 785 becomes m + (m + 77) = 785 can you solve this from here?

OpenStudy (therealmeeeee):

f + m = 785 m = 77 + f

OpenStudy (therealmeeeee):

subsstitue and solve with the above equations

OpenStudy (therealmeeeee):

77+ f + f = 785 77 + 2f = 785 2f = 708 f = 354

OpenStudy (therealmeeeee):

m = 77 + 354 m = 431

OpenStudy (therealmeeeee):

check 431 + 354 = 785 :)

OpenStudy (therealmeeeee):

and there u go do u understand now

OpenStudy (johnweldon1993):

But...as you can see....there are more FEMALES in the class...so there cannot be 431 males and 354 females....it would be switched...

OpenStudy (therealmeeeee):

yup

OpenStudy (therealmeeeee):

he not replying